OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. Answer y - 6 = -3x - 3 3x + y = 3 y = -3x + 3 y - 6 = -3(x + 1)

OpenStudy (mertsj):

y-6=-3(x+1) y-6=-3x-3